OpenStudy (anonymous):

Can anyone confirm this answer from this:1/3x+1/2y=7/3 1/3+3y=52/3 (-2,6)

OpenStudy (anonymous):

not true

OpenStudy (anonymous):

@cahit can you help me?

OpenStudy (anonymous):

Please anyone answer m, I really need help

OpenStudy (anonymous):

ok.. solve second equation.. then find the value of y (it is 17) , then put y in the first equation.. and find the value of x...

OpenStudy (anonymous):

3. Find the length of the slant height of a cone with a radius of 40 ft and a surface area of 9,420 square feet. Use 3.14 for pi. length = _____ ft 1

OpenStudy (anonymous):

17? how did you get that @cahit

OpenStudy (anonymous):

Can anyone answer me now

OpenStudy (anonymous):

Cahit, can you help me

OpenStudy (anonymous):

Please, anyone

OpenStudy (anonymous):

Cahit are you there

OpenStudy (anonymous):

\[1/3+3y=52/3\] \[3y=52/3-1/3\] \[3y=51/17\] \[3y=52/3\] \[3y=17\] \[y=17/3\]

OpenStudy (anonymous):

Cahit plz

OpenStudy (anonymous):

@cahit I am using elimination method, will that still work?

OpenStudy (anonymous):

there was a mistake in my solution the mistake is that 3y=51/17, erase it..

OpenStudy (anonymous):

@cahit are you not suppose to eliminate the fractions first?

OpenStudy (anonymous):

Anyone plz!!!!

OpenStudy (anonymous):

the easier sollution is substitution i think ..

OpenStudy (anonymous):

Cahit plz

OpenStudy (anonymous):

@cahit so we can use substitution then correct? What did you obtain from this?

OpenStudy (anonymous):

in elimination method you shuld eliminate y firstli, then find x and after that you can find y either in the first equation or in second equation..

OpenStudy (anonymous):

Yes my answer was -2,6

OpenStudy (anonymous):

ok.. then i must have any mistake ..

OpenStudy (anonymous):

Can you confirm that?

OpenStudy (anonymous):

@cahit CAn you confirm that my answer is correct?